• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z.-E. Meziani

    Log entry time 16:10:14 on July21,2005

    Entry number 148922

    keyword=Shift Summary

    8:00 Shift start.
    8:40 asked for beam off (target misbehaving)
    8:45 Controlled access 
    
    Jian Ping Chen to releive the target pressure
    David L'huillier to connect some cables for the Compton pol
    John Musson to get some IOC's working
    
    9:45 Hall A in Power power permit status
    
    Jian Ping changed the He bottle and set the pressure of the target around 200 
    psi. Temperature 6.4 K
    David reconnected the cables for the Compton signals to be available upstairs.
    John rebooted the IOC for the cavity monitors
    
    10:05 
    -Moved the target to empty position before asking for beam.
    Target didn't move first until have gone through the clear selected record.
    - Requested 5uA to check the beam position
    - Moving target to He position
    
    10:25 Starting with beam at 5 uA and ramping to 35 uA.
    There are still sometimes, for unclear reasons,  excursions of the target 
    pressure as large as +20/-10 psi 
    
    Starting production data
    
    Run #    Start --  Stop   # of events
    
    3738      11:00     11:51  91866 not useful (detectors saturated)
    
    Since the owl shift the high voltage on the detectors not set properly. Brian 
    Moffit adjusted the high voltages.
    
    Starting a new production run
    
    3740      12:12  12:18  11640 Short run 
    
    MCC took the beam back to tune up the Hall C beam line.
    12:40 Target moved to empty position
    12:45 Implementation of the Phase advance mode.
    Just starting to bring beam back to nominal BPMB and BPMA positions.
    13:05 MCC takes back beam for Hall C tuning
     
    13:50 Some beam is sent back to Hall A for tuning.
    Waiting for positions to settle to take look at the raster size with spot++.
    
    14:25
    A first look at the spot size, suggests that we need to adjust the size of the 
    raster. Beam off. Waiting for beam to adjust the raster.
    
    14:50
    Raster sizes with spot++ logged. The size at the target is only 3X3
    It is not adequate for production runs. We are requesting to load the old 
    optics in order for us to start production. See Paschke's log for the phase 
    advance effect.
    
    15:35 Beam is not back at the right position namely
    BPMA(1.6,0), BPMB(-1.4,0).
    
    
    15:35 Quench of the left septum magnet see the history charts of the septa and 
    the log. Marc Stevens needed to go down the Hall to recycle the power.
    16:00 Hall in controlled access end of shift